RPC weekly meeting

Europe/Zurich
5/R-026 (CERN)

5/R-026

CERN

15
Show room on map
Davide Boscherini (Universita e INFN, Bologna (IT)), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Mohamed Zaazoua (University of Science and Technology of China (CN))
Description

Weekly meeting about the RPC system activities

Zoom Meeting ID
61128581037
Host
Davide Boscherini
Useful links
Join via phone
Zoom URL
    • 09:00 09:15
      Communications to/from the gas group 15m
      Speakers: Beatrice Mandelli (CERN), Enrico Pastori (INFN e Universita Roma Tor Vergata (IT)), Gianluca Rigoletti (CERN), Giulio Aielli (INFN e Universita Roma Tor Vergata (IT)), Kacper Kapusniak (CERN), Roberto Guida (CERN)
    • 09:15 09:25
      RPC LVL1 status 10m
      Speakers: Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Gian Luigi Alberghi (Universita e INFN, Bologna (IT)), Marco Sessa (INFN e Universita Roma Tor Vergata (IT)), Massimo Della Pietra (University Federico II and INFN, Naples (IT)), Sabrina Perrella (Sapienza Universita e INFN, Roma I (IT))
    • 09:25 09:40
      RPC Dectector Status 15m
      Speakers: Daniela Katherinne Paredes Hernandez (University of Hong Kong (HK)), Eric Ballabene (University and INFN, Bologna (IT)), Mr Mohamed Zaazoua (University of Science and Technology of China (CN))
      • BML.C.11.CO.Ly0 HV channel driving high current on 29th July [e-log].
        • HV channel tripped after reaching ~250uA, no current visible in the gas gaps. HV channel ramped down to 9300V.

      • BML.B.09.PI.Ly0 HV channel driving high current on 29th July [e-log].
        • BML1A09.PI.DP2.Ly0.DZ1 gas gap driving high current. HV channel ramped down to 9300V.

      • BMS.A.04.CO.Ly0 HV channel driving high current on 29th July [e-log].
        • HV channel tripped after reaching ~180uA. Gas gap current are relatively not high. Icalib increased from 179uA to 185uA. HV channel ramped to nominal HV.

      • BML.B.15.CO.Ly1 HV channel driving high current on 29th July [e-log].
        •  HV channel tripped due to BML2C15.CO.DP2.Ly1.DZ1 gas gap driving high current. HV Vconfig set to 9200V.

      • BMS.A.16.CO.Ly0 HV channel driving high current on 30th July [e-log].
        • HV channel tripped due to BMS3A16.CO.Ly0.DZ2 and SPC4A16.CO.DP2.Ly0.DZ1 gas gaps driving high current. HV Vconfig set to 9200V.

      • BML.C.01.CO.Ly1 HV channel driving high current on 30th July [e-log].
        • HV channel tripped due to BML3C01.CO.DP1.Ly1.DZ2 and BML3C01.CO.DP2.Ly1.DZ2 gas gaps driving high current. HV channel ramped down to 9150V.

      • BML.C.11.CO.Ly0 HV channel driving high current on 30th July [e-log].
        • HV channel tripped due to BML3C01.CO.DP1.Ly1.DZ2 and BML3C01.CO.DP2.Ly1.DZ2 gas gaps driving high current. HV channel ramped down to 9150V.

      • RPC FATAL in side A/C and INFRASTRUCTURE due to 4 boards that went into unplugged status on 31st July [e-log, e-log, e-log].
        • Firstly masked, recovered the following day sending the reset command to the affected crates on the CAEN crate console. Recovered 4 A3512AP HV modules, 2 A3801 ADC modules. All mainframes were reset.

      • BOS.A2.02.CO.Ly1 HV channel driving high current on 1st August [e-log].
        • HV channel tripped due to BOS4A02.CO.DP2.Ly0.DZ1 gas gap on the other layer driving high current. HV channel ramped down to 8800V.

      • BOS.C2.16.CO.Ly1 HV channel driving high current on 2nd August [e-log].
        • Overcurrent already at 3660V, HV channel masked and turned off.
        • RPC resynchs in multiple sectors due to the noise bursts, leading to multiple stopless removals and towers killed [e-log].

      • BOL.C2.07.CO.Ly0 HV channel driving high current on 3rd August [e-log]. 
        • HV channel ramped down to 7000V due all connected gas gaps starting to drive high current from around 16:50.
    • 09:40 09:55
      Noise issue 15m
      Speakers: Claudio Luci (Sapienza Universita e INFN, Roma I (IT)), Mr Mohamed Zaazoua (University of Science and Technology of China (CN))
    • 09:55 10:05
      Resin update 10m
      Speakers: Ahmet Renklioglu (Istinye University (TR)), Dr Aydin Ozbey (Istanbul University-Cerrahpasa (TR))
      • A detailed brainstorming session was conducted with Elif Tug and Amine İclal Akın to address the issues that need to be resolved for R&D. Possible experimental setups for the systematic tests of the new injection machine were considered.

      • Our first objective is to test the quality of the foam produced when the machine is operated at a pressure that does not damage any components inside the box.

      • Additionally, the necessary arrangements were made for the training of two engineers/graduate students who will be working on Gas Repair Operations in the ATLAS Cavern throughout YETS 24/25. Their training in the laboratory is scheduled to be completed within three weeks.